# Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations

### v2

@@ 3 Problem Definition
This section introduces the core concepts and mathematical framework underlying our approach. We begin by defining the key notation and problem setup, followed by a formal description of how we quantify the impact of input perturbations. We then describe the use of a surrogate model to approximate model behavior locally and provide theoretical justification based on the Wasserstein distance [63, 112, 113]. Lastly, we outline a structured methodology for implementing a Wasserstein-based surrogate framework tailored to large language models [114].
@@ 3.1 Notation and Problem Setting
We aim to construct a local surrogate model that explains the behavior of a black-box large language model (LLM), denoted as $\pi^{(n)}$, in response to a natural language input prompt. The goal is to understand how specific parts of the input affect the model’s output distribution. The following notation is used:
@@ 3.2 Input-Level Distance
To quantify the semantic difference between the original prompt $x$ and each perturbed version $\hat{x}_{j}$, we compute the Word Mover’s Distance (WMD) [115]:
$\delta_{x_{j}}=\mathrm{WMD}(x,\hat{x}_{j}),$ (1)
@@ 3.3 Output-Level Distribution Shift
To assess the effect of input perturbation on the model’s behavior, we compute the Wasserstein distance between the output distributions of the original and perturbed inputs:
$\Delta(x,\hat{x}_{j})=W\left(\pi^{(n)}(y\mid x),\pi^{(n)}(y\mid\hat{x}_{j})\right),$ (2)
@@ 3.4 Weighting via Input Similarity
$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right),$ (3)
@@ 3.5 Fitting a Local Surrogate
where $\theta_{0}\in\mathbb{R}$ is a bias term and $\theta\in\mathbb{R}^{d}$ is the weight vector. Parameters $\theta$ are estimated by minimizing the weighted squared error.
@@
Where the loss ensures local fidelity between the surrogate and the observed output shifts.
@@ 3.6 Theoretical Justification (Lipschitz Smoothness)
We define a shift function that quantifies the model’s response to perturbation:
@@
If $\pi^{(n)}$ is Lipschitz continuous concerning its input, then the function $f$ is also Lipschitz:
@@
The notation $x_{a}$ and $\hat{x}_{a}$ represents a specific input and its corresponding perturbation. The subscript allows us to extend the formulation to multiple input pairs $(x_{i},\hat{x}_{i})$ when analyzing the Lipschitz continuity or stability across perturbations. $N_{pert}$ denotes the total number of perturbations.
[…]
@@
Where the total variation distance ($TV$) is given by:
@@
This bound highlights that small changes in distribution (in the TV sense) imply bounded Wasserstein distance.
@@ 3.7 Linear Surrogates Approximate Lipschitz Functions
3.8 Putting It All Together
The complete Wasserstein-based LIME surrogate procedure consists of:
  1. 1.
    Generating perturbations $\{\hat{x}_{j}\}_{j=1}^{J}$ near input prompt $x$.
  2. 2.
    Computing semantic distances $\delta_{x_{j}}$ (Equation 1) and output shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
  3. 3.
    Assigning similarity weights $w_{j}$ via Equation 3.
  4. 4.
    Extracting feature representations $z_{j}$ for each perturbation.
  5. 5.
    Fitting a local surrogate model $h_{\theta}$ by minimizing the loss (Equation 5).
  6. 6.
    Using $h_{\theta}$ to interpret the influence of each input feature on model behavior.
This framework offers a principled and interpretable approach to analyzing large language models through localized perturbation and response analysis.
@@ 4 PROPOSED METHOD
The proposed approach, named gSMILE (Generative SMILE), enhances the interpretability of large language models (LLMs) for text generation by analyzing how specific textual inputs influence the generated outputs. Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behavior [9, 114].
@@
Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as demonstrated in Fig. 5. Instead of perturbing the generated image, as in SMILE, we perturb the input prompt by including or excluding certain words. We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between each generated output and the original output [115]. This distance quantifies the difference in the production, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
These similarity scores are then used as the outcome variable in a weighted linear regression model, with weights derived from the text distances and perturbations [116]. This regression model helps determine how each word affects the generated text. The coefficients derived from this model quantify the influence of each word, which we use to create a visual heatmap highlighting the most influential words in the text prompt. This heatmap provides an intuitive visual representation of the elements of the prompt that contribute the most to the generated output.
Using this enhanced three-step process, which builds on SMILE’s methodology [11], we aim to help users understand how particular elements of text prompts impact text generation. This technique enhances user control and predictability in text generation models by providing insight into which aspects of textual input, such as stylistic instructions or descriptive terms, have the most significant impact on the final output.
The first step utilizes a text generation model that is trained on input text. The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output. The text generation model is expected to produce responses similar to the original response, varying only based on the critical components of the input texts [117].
Different text permutations are utilized from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into separate words. The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [118].
While the text generation model faces challenges in tasks such as counting objects and spatial reasoning, these issues can be mitigated by carefully selecting perturbation texts and restricting the input domain for perturbations [119]. The text generation models enable the proposed method to create responses that differ from the original text solely in terms of keywords related to the perturbation texts.
We use the Wasserstein distance to quantify these differences, comparing the distributions in both text embeddings [120, 112]. This metric is employed because it provides a more comprehensive measure of the differences between text embeddings compared to other ECDF-based distances, such as Kolmogorov–Smirnov or Cramér–von Mises. Although these methods focus on specific aspects, the Wasserstein distance captures both broad and subtle changes in the data [120, 113].
Finally, by applying linear regression, the space of input prompts is linked to the output space of the text generation model. As a result, the model’s behaviour can be interpreted by examining the regression coefficients [121, 116].
@@ 4.1 Text Generating and Perturbed Prompts
To begin, we generate variations of the original text prompt, referred to as “perturbed prompts,” to analyze how subtle changes in the input affect the output of a text generation model [122, 123]. The original prompt $p_{\text{org}}$ is first tokenized into individual words. From these tokens, multiple perturbed prompts $\{p_{\text{pert}}^{i}\}_{i=1}^{N}$ are constructed by selectively including or excluding specific words. Each perturbed prompt is then passed to the text generation model, which produces a corresponding output. The output generated from the original prompt, $t_{\text{org}}^{\prime}$, serves as the reference (or baseline) for comparison.
For the original text prompt $p_{\text{org}}$, the text generation model $\phi_{\text{gen}}$ produces the output:
$t_{\text{org}}^{\prime}=\phi_{\text{gen}}(p_{\text{org}})$ (11)
Similarly, for each perturbed prompt $p_{\text{pert}}^{i}\in S_{\text{pert}}$, the model generates:
$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in\{1,\dots,N\}$ (12)
Here, $\phi_{\text{gen}}$ is the text generation function, and $t_{\text{pert}}^{\prime i}$ denotes the generated output for the $i$-th perturbation. These outputs are used to measure how individual word-level changes in the prompt influence the model’s behavior [63, 118].
@@ 4.2 Creating the Interpretable Space
To represent the output space of the text generation model in a one-dimensional similarity space, we compute the Wasserstein distance between each generated text from a perturbed prompt and the output generated from the original prompt [112, 113]. Throughout this process, the input structure remains unchanged, and only the perturbed words differ across inputs.
@@
To capture the semantic similarity between generated outputs, we use Word Mover’s Distance (WMD) to embed the texts into a feature space [115]. These embeddings enable the focus to remain on high-level semantic content, thereby mitigating lexical noise and enhancing the reliability of the computed distances. The resulting scalar distances are used as the dependent variable in the subsequent regression model.
The Wasserstein distance between each perturbed output $t_{\text{pert}}^{\prime i}$ and the original output $t_{\text{org}}^{\prime}$ is computed as:
$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{P}\right)^{\frac{1}{P}},\quad\forall i\in\{1,\dots,N\}$ (13)
In this equation:
  • $\delta_{\text{WMD}}(\cdot)$ denotes the semantic embedding of a generated text,
  • $n$ is the dimensionality of the embedding space,
@@
  • $P$ specifies the norm order used in the Wasserstein metric,
  • and $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance computed for the $i$-th perturbation.
Using p-values is crucial for validating the significance of the observed Wasserstein distances (WD) between texts generated under different prompt perturbations. By calculating a p-value, we test whether the observed WD could have occurred by chance, determining if the perturbation-induced changes are meaningful [124]. To compute this p-value, we use the Bootstrap algorithm [125, 126], which estimates the probability of observing a WD as large as the observed one. This involves generating distributions of WDs through repeated sampling and comparing each sampled WD with the observed value. If a small proportion of sampled WDs exceed the observed WD, the resulting p-value confirms that the differences are statistically significant, supporting the validity of the perturbation effects.
@@
The Wasserstein-$P$ metric enables adjustment of the sensitivity of the Wasserstein Distance (WD) to different types of text variations by tuning the norm order $P$. Smaller values of $P$ emphasize fine-grained, local differences between distributions, while larger values capture broader, more global shifts [113]. By experimenting with various values of $P$, we can select the one that best captures the semantic differences introduced by prompt perturbations, thereby improving both the reliability and interpretability of the analysis [113].
To assess the statistical significance of the computed Wasserstein distances, we apply hypothesis testing using p-values. Distances associated with high p-values, indicating non-significant variation, are excluded from further consideration [124]. As detailed in Algorithm 1, the Bootstrap algorithm [127, 108] is used to estimate these p-values for each observed WD.
For example, let X and Y denote two sets of text-based embeddings. The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y. In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed. If boostWD exceeds the original observed WD, a counter (bigger) is incremented. After all iterations, the p-value (pVal) is estimated as the proportion of iterations where boostWD exceeds the observed WD, as outlined in Algorithm 1.
Input: Two sets of text-based distances: $X$ and $Y$
@@
Output: Observed Wasserstein Distance (WD) and corresponding p-value (pVal)
@@
[…]
6 $n\leftarrow\texttt{LX}+\texttt{LY}$;
@@ 7 $\texttt{bigger}\leftarrow 0$;
8
9for $i\leftarrow 1$ to MaxItr do
10       Draw a random sample $e$ of size LX from $XY$;
11       Draw a random sample $f$ of size LY from $XY$;
[…]
@@ 14             $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
[…]
Algorithm 1 Bootstrap Algorithm for Wasserstein Distance (WD) P-Value Calculation
Figure 5: gSMILE flowchart for Large Language Models
When generating perturbation texts from the original prompt, the key terms in the prompt are essential for maintaining semantic similarity between perturbed prompts and the source. To account for this, we define a similarity metric between each perturbed prompt and the original, which is then used as a sample weight in the interpretable model. This similarity is quantified using Word Mover’s Distance (WMD), which calculates the minimum cost to transport the embedded words of one document to match another [115].
The WMD between the original prompt $p_{\text{org}}$ and a perturbed prompt $p_{\text{pert}}^{i}$ is defined as:
$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\cdot d(w_{k},w_{l})$ (14)
subject to $\sum_{l=1}^{m}T_{kl}=d_{k},\quad\forall k=1,\dots,n$
$\sum_{k=1}^{n}T_{kl}=d^{\prime}_{l},\quad\forall l=1,\dots,m$
$T_{kl}\geq 0$
Here:
  • $w_{k}$ and $w_{l}$ denote the $k$-th and $l$-th words from the original and perturbed prompts respectively,
  • $d(w_{k},w_{l})$ is the distance (typically cosine or Euclidean) between the embeddings of $w_{k}$ and $w_{l}$,
  • $T_{kl}$ is the flow variable indicating how much of $w_{k}$’s weight is transported to $w_{l}$,
  • $d_{k}$ and $d^{\prime}_{l}$ represent the normalized bag-of-words weights for the original and perturbed texts.
To use this distance as a weighting factor in the regression model, we normalize it using a Gaussian kernel:
$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})=\exp\left(-\left(\frac{\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma}\right)^{2}\right),\quad\forall i\in\{1,\dots,N\}$ (15)
Here, $\sigma$ is a kernel width parameter that controls sensitivity to dissimilar inputs. The resulting similarity score $\pi_{i}$ falls in the range $[0,1]$, assigning higher weights to perturbed prompts that are semantically closer to the original.

### v4

@@ 3 Problem Definition
This section introduces the core concepts and mathematical framework underlying our approach. We begin by defining the key notation and problem setup, followed by a formal description of how we quantify the impact of input perturbations. We then describe the use of a surrogate model to approximate model behaviour locally and provide theoretical justification based on the Wasserstein distance [64, 105, 106]. Lastly, we outline a structured methodology for implementing a Wasserstein-based surrogate framework tailored to large language models [62].
Word Mover’s Distance (WMD) is the 1-Wasserstein (Earth Mover’s) distance between two documents represented as normalised bags-of-words in an embedding space, using the Euclidean distance between word embeddings as the ground cost [107]. WMD provides a semantically meaningful measure of dissimilarity between documents by capturing the minimum amount of semantic “movement” needed to align their word distributions, making it well-suited for our interpretability framework.
@@ 3.1 Notation and Problem Setting
We aim to construct a local surrogate model that explains the behaviour of a black-box large language model (LLM), denoted as $\pi^{(n)}$, in response to a natural language input prompt. The goal is to understand how specific parts of the input affect the model’s output distribution. The following notation is used:
@@ 3.2 Input-Level Distance
To quantify the semantic difference between the original prompt $x$ and each perturbed version $\hat{x}_{j}$, we compute the Word Mover’s Distance (WMD) [107]. We denote this input-level semantic distance as IWMD (Inner WMD):
$\delta_{x_{j}}=\mathrm{IWMD}(x,\hat{x}_{j}),$ (1)
@@ 3.3 Output-Level Distribution Shift
To assess the effect of input perturbation on the model’s behaviour, we compute the Wasserstein distance between the output distributions of the original and perturbed inputs. We refer to this as OWMD (Outcome WMD):
$\mathrm{OWMD}(x,\hat{x}_{j})=W\left(\pi^{(n)}(y\mid x),\pi^{(n)}(y\mid\hat{x}_{j})\right),$ (2)
$\Delta(x,\hat{x}_{j})=\mathrm{OWMD}(x,\hat{x}_{j}),$ (3)
@@ 3.4 Weighting via Input Similarity
$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma}\right)^{2}\right),$ (4)
@@ 3.5 Fitting a Local Surrogate
where $\theta_{0}\in\mathbb{R}$ is a bias term and $\theta\in\mathbb{R}^{d}$ is the weight vector. Parameters $\theta$ are estimated by minimising the weighted squared error.
@@
Where the loss ensures local ATT fidelity between the surrogate and the observed output shifts.
@@ 3.6 Theoretical Justification (Lipschitz Smoothness)
We define a shift function that locally quantifies the model’s response to perturbation:
@@
If $\pi^{(n)}$ is Lipschitz continuous with respect to its input, then the function $f$ is also Lipschitz [108]:
@@
The notation $x_{a}$ and $\hat{x}_{a}$ represents a specific input and its corresponding perturbation. The subscript allows us to extend the formulation to multiple input pairs $(x_{i},\hat{x}_{i})$ when analysing the Lipschitz continuity or stability across perturbations. $N_{pert}$ denotes the total number of perturbations.
where $L>0$ is the Lipschitz constant, and $\|\cdot\|$ denotes a norm over the input space (e.g., token embedding space).
@@
Where $\text{diam}(\mathcal{Y})=\sup_{y_{1},y_{2}\in\mathcal{Y}}d(y_{1},y_{2})$ denotes the diameter of the output space under the chosen ground metric $d$, i.e., the maximum possible distance between any two outputs.
The total variation distance ($TV$) is given by:
@@
This bound highlights that small changes in distribution (in the TV sense) imply a bounded Wasserstein distance.
@@ 3.7 Linear Surrogates Approximate Lipschitz Functions
The complete Wasserstein-based LIME surrogate procedure consists of generating perturbations $\{\hat{x}_{j}\}_{j=1}^{J}$ near input prompt $x$. The process then computes semantic distances $\delta_{x_{j}}$ (Equation 1) and output shifts $\Delta(x,\hat{x}_{j})$ (Equation 3). Following this, similarity weights $w_{j}$ are assigned via Equation 4. The procedure continues by extracting feature representations $z_{j}$ for each perturbation and fitting a local surrogate model $h_{\theta}$ by minimising the loss (Equation 6). Finally, the fitted model $h_{\theta}$ is used to interpret the influence of each input feature on model behaviour. This framework offers a principled and interpretable approach to analysing large language models through localised perturbation and response analysis.
@@ 4 Proposed Method
The proposed approach, named gSMILE (Generative SMILE), enhances the interpretability of large language models (LLMs) for text generation by analysing how specific textual inputs influence the generated outputs. Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behavior [109, 62].
@@
Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as illustrated in Fig. 5. Instead of perturbing the generated image, as in the original SMILE framework, we perturb the input prompt. Each perturbed prompt $\hat{x}_{j}$ produces a corresponding text output, from which we compute the outcome-level semantic shift $\Delta(x,\hat{x}_{j})$ using the IWMD as defined  1. This allows us to identify which words in the input prompt most influence the model’s output.
The set of output shifts $\{\Delta(x,\hat{x}_{j})\}_{j=1}^{J}$ is then used as the outcome variable in a weighted linear regression model, with weights $w_{j}$ computed from the input-level semantic distances $\delta_{x_{j}}$ via a Gaussian kernel (Eq. 4). This regression estimates how each input feature (token) influences the generated output within the local neighbourhood of $x$. The resulting coefficients $\theta$ quantify the relative importance of each token, which we visualise as a heatmap highlighting the most influential words in the prompt.
@@
Using this enhanced three-step process, which builds on SMILE’s methodology [13], we aim to help users understand how particular elements of text prompts impact text generation. This approach improves user control and predictability by showing which textual components, such as stylistic instructions or descriptive terms, most influence the final output.
In the first step, we employ a black-box text generation model $\pi^{(n)}$ that receives an original prompt $x$ and produces an associated output $\pi^{(n)}(y\mid x)$. We then construct a set of perturbed prompts $\{\hat{x}_{j}\}_{j=1}^{J}$ by applying small, semantically meaningful edits to $x$ (e.g., removing or replacing specific tokens). For each $\hat{x}_{j}$, the model generates $\pi^{(n)}(y\mid\hat{x}_{j})$. Ideally, these outputs remain similar to the original output, with variation arising primarily from the altered components of the prompt [110].
We quantify the distributional shift between the original and perturbed outputs using the OWMD as defined 2 $\Delta(x,\hat{x}_{j})$, as defined in Eq. 3, computed in the output embedding space [111, 105]. This metric is chosen for its ability to capture both broad and subtle differences compared to ECDF-based measures [106].
Finally, we fit a weighted linear surrogate model $h_{\theta}$ (Eq. 5) by minimising the loss in Eq. 6, linking the feature representation $z_{j}$ of each perturbed prompt to its corresponding $\Delta(x,\hat{x}_{j})$. The learned coefficients $\theta$ provide an interpretable mapping between input prompt components and changes in the model’s generated output [112, 113].
Figure 5: gSMILE flowchart for Large Language Models
@@ 4.1 Text Generating and Perturbed Prompts
We begin by generating variations of the original text prompt $x$, referred to as “perturbed prompts” $\{\hat{x}_{j}\}_{j=1}^{J}$, to analyse how subtle changes in the input affect the output of a text generation model [114]. The original prompt $x$ is first tokenised into individual tokens. From these tokens, multiple perturbed prompts are constructed by selectively including, excluding, or substituting specific words or phrases. Each perturbed prompt $\hat{x}_{j}$ is then passed to the text generation model $\pi^{(n)}$, which produces a corresponding output. The output generated from the original prompt, $\pi^{(n)}(y\mid x)$, serves as the reference (or baseline) for comparison.
Formally, the outputs for the original and perturbed prompts follow the notation in Section 3:
the model generates $y_{\text{org}}\sim\pi^{(n)}(y\mid x)$ for the original prompt, and $y_{\text{pert},j}\sim\pi^{(n)}(y\mid\hat{x}_{j})$ for each perturbation $j\in\{1,\dots,J\}$.
Here, $\pi^{(n)}$ denotes the black-box text generation model, $y_{\text{org}}$ is the output for the original prompt, and $y_{\text{pert},j}$ denotes the output for the $j$-th perturbation. These outputs are subsequently used in the computation of the output shift $\Delta(x,\hat{x}_{j})$ (Eq. 3) to measure how individual token-level changes in the prompt influence the model’s behavior [64, 115].
@@ 4.2 Creating the Interpretable Space
To represent the output space of the text generation model in a one-dimensional similarity space, we compute the OWMD distance between the output for each perturbed prompt $\hat{x}_{j}$ and the output for the original prompt $x$ (as defined in Eq. 3[105, 106]. Throughout this process, the input structure remains unchanged, and only the perturbed tokens differ across inputs.
@@
To capture semantic similarity between generated outputs, we employ the Outcome Word Mover’s Distance (OWMD) to measure distances between texts based on their representations in a semantic embedding space [107]. Operating in this embedding space focuses the comparison on high-level semantic content, thereby mitigating lexical noise and improving the reliability of the computed distances. The resulting scalar values $\Delta(x,\hat{x}_{j})$ are then used as the dependent variable in the weighted regression model.
@@
Here, $\phi(\cdot)$ denotes the output embedding function, $n$ is the dimensionality of the embedding space, and $P$ specifies the norm order used in the Wasserstein metric.
For each perturbation $j$, $\Delta(x,\hat{x}_{j})$ measures the distributional shift in the generated output relative to the baseline.
After computing $\Delta(x,\hat{x}_{j})$ for each perturbation, we assess their statistical significance using p-values [116]. This ensures that only meaningful output shifts are considered in the regression model. A p-value quantifies the probability that an observed Outcome Word Mover’s Distance (OWMD) between outputs could have occurred purely by random chance, under the null hypothesis that the perturbation has no real effect.
@@
To control the sensitivity of OWMD to different types of changes, we use the Wasserstein-$P$ metric [106]. The norm order $P$ determines the scale of differences captured: smaller $P$ values emphasise fine-grained, local variations, whereas larger $P$ values capture broader, global shifts. By experimenting with different $P$ values, we select the one that best captures the semantic effects of prompt perturbations, thereby improving both the robustness and interpretability of the analysis.
We compute p-values for each observed OWMD using a Bootstrap resampling procedure [117, 104], as detailed in Algorithm 1. This non-parametric approach repeatedly resamples the combined set of embeddings from both the original and perturbed outputs to construct an empirical distribution of OWMD values under the null hypothesis. The p-value is then estimated as the proportion of bootstrap distances greater than the observed OWMD. Perturbations with high p-values, indicating non-significant differences, are excluded from further interpretation.
Input: Two sets of text embeddings: $X$ and $Y$
@@
Output: Observed Outcome Wasserstein Distance (OWMD) and p-value (pVal)
@@
3 $\texttt{OWMD}\leftarrow\texttt{Wasserstein\_Dist}(X,Y)$;
4 $\texttt{XY}\leftarrow X\cup Y$;
5 $L_{X}\leftarrow|X|$, $L_{Y}\leftarrow|Y|$;
@@ 6 $\texttt{bigger}\leftarrow 0$;
7
8for $i\leftarrow 1$ to MaxItr do
9   Sample $e$ of size $L_{X}$ from XY without replacement;
10   Sample $f$ of size $L_{Y}$ from XY without replacement;
11 $\texttt{bootOWMD}\leftarrow\texttt{Wasserstein\_Dist}(e,f)$;
12 if $\texttt{bootOWMD}>\texttt{OWMD}$ then
@@ 13    $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
19 return $\texttt{OWMD},\texttt{pVal}$;
Algorithm 1 Bootstrap-based P-Value Calculation for Outcome Wasserstein Distance (OWMD)
When generating perturbed prompts from the original prompt $x$, preserving key terms is essential for maintaining semantic similarity between each perturbed prompt $\hat{x}_{j}$ and the source. To capture this similarity, we compute the input-level semantic distance $\delta_{x_{j}}$ using the Inner Wasserstein Distance over tokens (IWMD) [107], as introduced in Eq. 1. This value is later transformed into a sample weight $w_{j}$ for the interpretable surrogate model (Eq. 4).
To incorporate input similarity into the regression model, each IWMD value is converted into a similarity weight via a Gaussian kernel, following Eq. 4:
$w_{j}(x,\hat{x}_{j})=\exp\left[-\left(\frac{\mathrm{IWMD}(x,\hat{x}_{j})}{\sigma}\right)^{2}\right],\quad\forall j\in\{1,\dots,J\}.$ (12)
Here, $\sigma>0$ is the kernel width parameter that controls the rate of decay in weight as semantic dissimilarity increases. The resulting weights $w_{j}\in(0,1]$ assign greater importance to perturbed prompts that are more semantically similar to the original prompt $x$, and down-weight those that differ substantially.